If the smaller of two numbers is one-half of the larger number and the sum of the two numbers is 57, find the numbers.
Answer by ikleyn(52876)

.

    smaller + larger = 57

       x    +   2x   = 57

           3x        = 57

            x        = 57/3 = 19   the smaller.


ANSWER.  The smaller number is 19.  The greater number is 2*19 = 38.

Solved.